OpenStudy (anonymous):

how do you solve this problem? " two combines harvest a field in 10 hours, if one combine can complete the job in 30 hours, how long should it take the other combine to harvest the field alone? "

OpenStudy (anonymous):

1 C completes 1 Field in 30 hrs .... > in ! hr 1 C completes 1/30 Field 2 C complete 1 Field in 10 hrs ...> In 1 hr 2 C complete 1/10 Field ... > in 1 hr the second C completes ( 1/10 - 1/30) = 2/30 =1/15 Field ..> he completes 1 field in 15 hrs

OpenStudy (anonymous):

let us suppose their are two people A & B A can complete the job in 30 hrs so in one hour he will complete (1/30) th job simmilarly B can complete the job in one hour = (1/10) th job therefore in a hour both will do = (1/30)+(1/10) there fore both together will complete the work in (15/2) hrs = 7.5 hrs

OpenStudy (anonymous):

i was telling u the process to calculate

OpenStudy (anonymous):

they complete the field in 10 when together

OpenStudy (anonymous):

sagar i was telling the process used so i had assumed two different peersons A and B

OpenStudy (anonymous):

so is it 15 hours or 7.5? I'm confused. o:

OpenStudy (anonymous):

hey i was telling u the procedure in this case your answer is 15 hrs. i was just explaining u the things how to calculate

OpenStudy (anonymous):

oh okay, thank you.